OpenStudy (anonymous):

Simplify: log2(1024n/log2n)

OpenStudy (anonymous):

Not sure what property would help here.

ganeshie8 (ganeshie8):

\(\large \mathbb{\frac{\log_2 1024^n}{\log_2 n}}\)

ganeshie8 (ganeshie8):

is the question like that ?

OpenStudy (anonymous):

the bottom part is log2(4sqrt(8))

ganeshie8 (ganeshie8):

\(\large \mathbb{\frac{\log_2 1024^n}{\log_2 \sqrt{8}}}\)

ganeshie8 (ganeshie8):

like that ? is the top part correct ?

OpenStudy (anonymous):

Its log2(1024n/log2n) / log2(4sqrt(8)) the bottom being quad root of 8

OpenStudy (jdoe0001):

\(\bf \large \cfrac{log_2\left(\frac{1024^n}{log_2n}\right)}{log_2\sqrt[4]{8}}\quad ?\)

OpenStudy (anonymous):

Yes, that is correct

OpenStudy (anonymous):

execpt it is 1024n, no exponent

OpenStudy (jdoe0001):

ok

OpenStudy (jdoe0001):

http://www.chilimath.com/algebra/advanced/log/images/rules%20of%20exponents.gif <--- notice those rules \(\bf \large{ {\color{blue}{ 1024\implies 2^{10}\qquad 8\implies 2^3\qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}=a^{\frac{{\color{blue} n}}{{\color{red} m}}}}} \\ \quad \\ \quad \\ \cfrac{log_2\left(\frac{1024n}{log_2n}\right)}{log_2\sqrt[4]{8}}\implies \cfrac{log_2\left(\frac{2^{10}n}{log_2n}\right)}{log_2\sqrt[4]{2^3}}\implies \\ \quad \\ \cfrac{log_2\left(\frac{2^{10}n}{log_2n}\right)}{log_2(2^{\frac{3}{8}})}\implies \cfrac{log_2(2^{10}n)-log_2n}{log_2(2^{\frac{3}{8}})}}\) see if you can apply the 1st log rule listed on the picture to the numerator

OpenStudy (jdoe0001):

hmmm I have an 8 that came out of nowhere....lemme fix that \(\bf \large{ {\color{blue}{ 1024\implies 2^{10}\qquad 8\implies 2^3\qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}=a^{\frac{{\color{blue} n}}{{\color{red} m}}}}} \\ \quad \\ \quad \\ \cfrac{log_2\left(\frac{1024n}{log_2n}\right)}{log_2\sqrt[4]{8}}\implies \cfrac{log_2\left(\frac{2^{10}n}{log_2n}\right)}{log_2\sqrt[4]{2^3}}\implies \\ \quad \\ \quad \\ \cfrac{log_2\left(\frac{2^{10}n}{log_2n}\right)}{log_2(2^{\frac{3}{4}})}\implies \cfrac{log_2(2^{10}n)-log_2n}{log_2(2^{\frac{3}{4}})}}\)

OpenStudy (jdoe0001):

say what would you get if you were to apply the 1st rule listed at http://www.chilimath.com/algebra/advanced/log/images/rules%20of%20exponents.gif to say \(\bf log_2(2^{10}n)\quad ?\)